OpenStudy (anonymous):

The 1st electron in the H2 atom is ejected from the atom after absorbing a total amt of Energy = 20.4eV . Calculate its de Broglie wavelenght in A^0

OpenStudy (experimentx):

what's the first ionization potential for He?

OpenStudy (anonymous):

E = 2.18 * 10^-18 * 2

OpenStudy (anonymous):

i think...

OpenStudy (anonymous):

but the question asks abt hydrogen

OpenStudy (anonymous):

E = 13.62 eV ------ Hydrogen

OpenStudy (experimentx):

oh ... i thought i say He ... i think i should get some proper sleep.

OpenStudy (experimentx):

do the usual way ... \[ \Delta E = 20 - 13.6 = 6.4eV \\ \lambda = {h \over p} = {h \over \sqrt{ 2 E }}\]

OpenStudy (anonymous):

we have to change eV to jolue.....right

OpenStudy (experimentx):

looks like i missed something ... i should really get some sleep. yes you should